State with a reason, whether each of the following shows oxidation or reduction a)1²(Aq)+2e -----------21(Aq)

B)cu(AQ)-------------------cu²(aq)+e

C)2Br(aq)-----------------br²(aq+2e)


1
Expert's answer
2022-01-17T14:55:04-0500

Solution:

a)

I2(s) + 2e → 2I(aq) (gain electrons = reduction half reaction)

The oxidation state of I changes from 0 to −1. Hence, I2 is reduced.


b)

Cu(s) → Cu2+(aq) + 2e

Cu(s) − 2e → Cu2+(aq) (lose electrons = oxidation half reaction)

The oxidation state of Cu changes from 0 to +2. Hence, Cu is oxidized.


c)

2Br(aq) → Br2(l) + 2e

2Br(aq) − 2e → Br2(l) (lose electrons = oxidation half reaction)

The oxidation state of Br changes from −1 to 0. Hence, Br ion is oxidized.